OpenStudy (anonymous):

What is the solution of the system below? x = -3y - 4 -3y = 2x - 7 z = -6x +y The possible Answers are: x=-5,y=11,z=-71 x=11,y=-71,z=-5 x=-71,y=-5,z=11 x=11,y=-5,z=-71

OpenStudy (luigi0210):

Have you made an attempt?

OpenStudy (anonymous):

Yes. I came up with the last one

OpenStudy (luigi0210):

Yup, you're right.

OpenStudy (anonymous):

Thanks, I needed the reassurance :)

OpenStudy (luigi0210):

You're welocme!

OpenStudy (luigi0210):

*welcome

OpenStudy (sleepyjess):

x = -3y - 4 -3y = 2x - 7 -3y = 2(-3y - 4) - 7 -3y = -6y - 8 - 7 -3y = -6y - 15 -3y + 6y = -6y + 6y - 15 3y = -15 y = -5 _____ -3y = 2x - 7 -3(-5) = 2x - 7 15 = 2x - 7 15 + 7 = 2x 2x = 22 x = 11 _____ z = -6x +y z = -6(11) + (-5) z = -66 - 5 z = -71